OpenStudy (anonymous):

Find the solution of the equation: 2ln(2+x)=1

OpenStudy (anonymous):

First divide both sides by 2

OpenStudy (anonymous):

x=e1/2-2

OpenStudy (anonymous):

i got x=-3 x=-1

OpenStudy (anonymous):

This leaves you with ln(2 + x) = 1/2 Recall that (ln) = (log base e) With that said we can rewrite the equation, using the logarithm laws. 2 + x = e^(1/2) subtract 2 from both sides x = e^(1/2) - 2

OpenStudy (anonymous):

what i did was ln(2+x)^2=1 x^2+4x+4=1

OpenStudy (anonymous):

that is wrong

OpenStudy (anonymous):

If you wanted to solve it that way it would be ln(x + 2)^2 = 1 (x + 2)^2 = e x^2+4x+4 = e x^2+4x+4 - e = 0 then solve it using quadratic formula or something and you will get your answer. But this will be a very messy process

OpenStudy (anonymous):

in your first comment do you mean multiply by 2 or 1/2

OpenStudy (anonymous):

I mean divide both sides by 2, which is same as saying multiply both sides by 1/2

OpenStudy (anonymous):

oh okay thanks a lot